# On Weak Compactness and Uniform Regularity in Lipschitz free Spaces

### v1

@@ 2. $(b)$
Let $M$ be a complete pointed metric space. For every $n\in\mathbb{N}$, let $\gamma_{n}=\sum_{k}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}\in S_{\mathcal{F}(M)}$ be convex sum/series of molecules such that $\sum_{k}d(x_{k}^{n},y_{k}^{n})\to 0$ as $n\to\infty$. Then $(\gamma_{n})_{n}$ has a subsequence which is an (asymptotically isometric) $\ell_{1}$-basis.
@@ 1.2. Lipschitz free spaces
Then $\mathsf{M}_{h}$ is a weak-to-weak continuous linear operator from $\operatorname{Lip}_{0}(K)$ into $\operatorname{Lip}_{0}(M)$, called a multiplication operator. Its preadjoint $T_{h}:\mathcal{F}(M)\to\mathcal{F}(K)$, given by $T_{h}(\delta(x))=h(x)\,\delta(f(x))$, is referred to as a weighted Lipschitz operator. Precise conditions on $h$ and $f$ ensuring boundedness of $\mathsf{M}_{h}$ and $T_{h}$ are characterized in [2].
@@ 3. $(iii)$
Let $\gamma_{1}\in W$ be such that $\|\gamma_{1}\|>\varepsilon$. Let $\mu_{1}\in\mathcal{F}(B(0,r))$ be such that $\|\gamma_{1}-\mu_{1}\|\leq\frac{\varepsilon}{4}$. Then $\|\mu\|>\frac{3\varepsilon}{4}$ and thus there exists $g_{1}\in B_{\operatorname{Lip}_{0}(M)}$ such that $\langle g_{1},\mu_{1}\rangle>\frac{3\varepsilon}{4}$. We may additionally assume that $g_{1}\mathord{\upharpoonright}_{B(0,2r)^{c}}=0$ by replacing $g_{1}$ with a function $g^{\prime}_{1}$ defined by $g^{\prime}_{1}=g_{1}$ on $B(0,r)$, $g^{\prime}_{1}=0$ on $B(0,2r)^{c}$, and extended to $M$ by McShane-Whitney theorem (see e.g. [24, Theorem 1.33]). Furthermore, $\langle g_{1},\gamma_{1}\rangle>\frac{\varepsilon}{2}$. Pick a finitely supported element $\gamma_{1}^{\prime}\in\mathcal{F}(M)$ such that $\langle g_{1},\gamma_{1}^{\prime}\rangle>\frac{\varepsilon}{2}$ and $\|\gamma_{1}-\gamma_{1}^{\prime}\|<\frac{\varepsilon\delta}{16r}$. Let $F_{1}=\operatorname{supp}(\gamma_{1}^{\prime})\cup\left\{{0}\right\}$ and define $f_{1}\colon[F_{1}]_{\delta/2}^{c}\cup F_{1}\rightarrow\mathbb{R}$ in such way that $f_{1}\mathord{\upharpoonright}_{[F_{1}]_{\delta/2}^{c}}=0$ and $f_{1}\mathord{\upharpoonright}_{F_{1}}=\frac{\delta}{4r}g_{1}\mathord{\upharpoonright}_{F_{1}}$. Since $g_{1}\mathord{\upharpoonright}_{B(0,2r)^{c}}=0$, for every $u\in F_{1}$ we have $\left|{g_{1}(u)}\right|\leq 2r$ and thus $\left|{f_{1}(u)}\right|\leq\frac{\delta}{2}$. Thus, since $\frac{\delta}{4r}\leqslant 1$, the Lipschitz constant of $f_{1}$ is at most 1. Extend $f_{1}$ to $M$ using McShane–Whitney Theorem. Then $f_{1}\in B_{\operatorname{Lip}_{0}(M)}$ and $\langle f_{1},\gamma_{1}^{\prime}\rangle=\frac{\delta}{4r}\langle g_{1},\gamma_{1}^{\prime}\rangle>\frac{\varepsilon\delta}{8r}$, and thus $\langle f_{1},\gamma_{1}\rangle>\frac{\varepsilon\delta}{16r}$.
@@ 3. Uniform regularity versus weak compactness
Note that for every $i\in\{1,\ldots,n\}\setminus I$, there exist $x_{i}\in B(u_{i},r_{i})\cap U^{c}$ and $y_{i}\in B(v_{i},r_{i})\cap U^{c}$. Thus, by [22, Lemma 1.2]:
@@
for every $k\in\mathbb{Z}$. The well-known Kalton decomposition [20, Proposition 4.3] asserts that there exist Lipschitz functions $\varphi_{n}:M\to\mathbb{R}$ with $\operatorname{supp}(\varphi_{n})\subset M_{n}$ such that the pre-adjoints $T_{n}=T_{\varphi_{n}}$ of the associated multiplication operators $\mathsf{M}_{\varphi_{n}}:\operatorname{Lip}_{0}(M_{n})\to\operatorname{Lip}_{0}(M)$ satisfy that
@@ 2. $(b)$
Notice that when $\mu$ is finite, condition (a) implies condition (b). According to the Dunford–Pettis theorem, a set $H$ is relatively weakly compact in $L^{1}(\mu)$ if and only if it is equi-integrable (see e.g. [23, Theorem 15.4]). It is worth noting that, since $L^{1}$-spaces are weakly sequentially complete (see [3, Theorem 5.2.9]), weakly precompact sets coincide with relatively weakly compact sets. In what follows, we will transfer the equi-integrability conditions through the isometry $L^{1}(\lambda_{T})\to\mathcal{F}(T)$ and describe the resulting properties in $\mathcal{F}(T)$.
@@ 5. Sequences of sums of “small molecules” contain an $\ell_{1}$-basis
The previous section highlights the significance of approximation by convex series of molecules. Informally, within a weakly compact set, the weight of molecules formed by pairs of points in close proximity should be controlled in a uniform way. In this section, we establish that this condition is, in a sense, necessary. Specifically, consider a sequence of convex series of molecules, denoted by $\gamma_{n}=\sum_{k}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}$, satisfying $\sum_{k}d(x_{k}^{n},y_{k}^{n})\to 0$ as $n\to\infty$. We will demonstrate that this sequence contains an $\ell_{1}$-basis, thereby proving that such a sequence cannot be contained in a weakly precompact set.
We shall start with the following lemma, whose proof relies on a well known fact of optimal transport theory that can be found e.g. in [11, Theorem 2.4]. For convenience of the reader, we recall the statement of this result. Let $(x_{n},y_{n})$ be pairs of distinct points in $M$. Then the following are equivalent:
  • The expression $\sum_{n}\lambda_{n}m_{x_{n}y_{n}}$ is a convex sum/series of molecules for some choice of $\lambda_{n}>0$ (hence any choice of $\lambda_{n}\geqslant 0$) with $\sum_{n}\lambda_{n}<\infty$.
  • There exists $f\in S_{\operatorname{Lip}_{0}(M)}$ such that $\left<{f,m_{x_{n}y_{n}}}\right>=1$ for all $n$.
  • The pairs $(x_{n},y_{n})$ are cyclically monotonic, that is: for every finite sequence of indices $i_{1},\ldots,i_{m}$, we have
    $d(x_{i_{1}},y_{i_{1}})+d(x_{i_{2}},y_{i_{2}})+\ldots+d(x_{i_{m}},y_{i_{m}})\leqslant d(x_{i_{1}},y_{i_{2}})+d(x_{i_{2}},y_{i_{3}})+\ldots+d(x_{i_{m}},y_{i_{1}}).$
@@
Let $M$ be a pointed metric space. Let $\mu:=\sum_{k}\lambda_{k}m_{x_{k}y_{k}}\in\mathcal{F}(M)$ be a convex sum/series of molecules and $\gamma\in\mathcal{F}(M)$. Let $\delta:=\sum_{k}d(x_{k},y_{k})$ and
@@
By an obvious approximation argument, we may assume that $\gamma$ and $\mu$ are finitely supported, with $\mu$ written as $\mu:=\sum_{k=1}^{n}\lambda_{k}m_{x_{k}y_{k}}$. Let $A=\operatorname{supp}(\gamma)\cup\left\{{0}\right\}$ and $B=\left\{{x_{1},y_{1},\ldots,x_{n},y_{n}}\right\}$. We may further assume that $\gamma$ is written as a convex sum of molecules from points in $A$, say $m_{x_{n+1}y_{n+1}},\ldots,m_{x_{N}y_{N}}$, where $x_{n+1},y_{n+1},\ldots,x_{N},y_{N}\in A$. Finally, we assume without loss of generality that $M=A\cup B$. Note that $A\cap B=\emptyset$ as $d(A,B)=C>\delta>0$.
@@
Claim: The pairs $(x_{k},y_{k})$ for $k=1,\ldots,N$ are cyclically monotonic with respect to $b$.
Let us assume that the claim is proved. By [11, Theorem 2.4], there exists $f\in S_{\operatorname{Lip}_{0}(M,b)}$ which norms all molecules making up $\gamma$ and $\mu$ in $\mathcal{F}(M,b)$. By definition of $b$, $f$ also norms all of those molecules in $\mathcal{F}(M,d)$. Since $f$ is $(1+\delta/C)$-Lipschitz with respect to $d$, one has
$\|\mu\|+\|\gamma\|=\left<{f,\mu+\gamma}\right>\leqslant\left\|{f}\right\|\left\|{\mu+\gamma}\right\|\leqslant\left({1+\frac{\delta}{C}}\right)\left\|{\mu+\gamma}\right\|.$
To conclude the proof, it only remains to prove the claim. Note that, since $\mu$ is a convex sum of the molecules $m_{x_{1}y_{1}},\ldots,m_{x_{n}y_{n}}$, and $\gamma$ is a convex sum of the molecules $m_{x_{n+1}y_{n+1}},\dots,m_{x_{N}y_{N}}$, by [11, Theorem 2.4] the sequences $(x_{k},y_{k})_{k=1}^{n}$ and $(x_{k},y_{k})_{k=n+1}^{N}$ are cyclically monotonic with respect to $d$, hence also with respect to $b$. Now let us consider $i_{1},\ldots,i_{m}\in\{1,\ldots,N\}$, with $m\geqslant 2$ and indices from both sets $\{1,\ldots,n\}$ and $\{n+1,\ldots,N\}$. Note that we may assume that the indices $i_{1},\ldots,i_{m}$ are pairwise distinct. Let $j_{1}<\ldots<j_{l}$ be such that the indices $i_{j_{k}}$, $1\leqslant k\leqslant l$, are those which belong to the set $\left\{{n+1,\ldots,N}\right\}$, and let us write $I=\{j_{1},\ldots,j_{l}\}$. To simplify the notation in the next estimates, we let $j_{l+1}:=j_{1}$, $i_{m+1}:=i_{1}$. Using the cyclic monotonicity of $(x_{k},y_{k})_{k=n+1}^{N}$ and the definition of $\delta$, we first obtain:
@@
Let $M$ be a pointed metric space. Let $\mu:=\sum_{k}\lambda_{k}m_{x_{k}y_{k}}\in\mathcal{F}(M)$ be a convex sum/series of molecules and $\gamma\in\mathcal{F}(M)$. Let $r>0$ be such that for every $k$ there exists $u\in\operatorname{supp}(\gamma)\cup\left\{{0}\right\}$ such that
@@
Using the two previous lemmas, we obtain the following generalization of [19, Theorem 2.6], which covers the case where each $\gamma_{n}$ is a molecule.
@@
Let $(\gamma_{n})$ be a sequence in $S_{\mathcal{F}({M})}$ such that each $\gamma_{n}$ is a convex sum/series of molecules expressed as follows:
$\gamma_{n}=\sum_{k}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}},\;\text{ with }\;\sum_{k}\left|{\lambda_{k}^{n}}\right|=\|\gamma_{n}\|=1.$
If $\lim_{n\to\infty}\sum_{k}d(x_{k}^{n},y_{k}^{n})=0,$ then $\lim_{n}\left\|{\gamma+\gamma_{n}}\right\|=\left\|{\gamma}\right\|+1$ for all $\gamma\in\mathcal{F}({M})$.

### v2

@@ 2. $(b)$
Let $M$ be a complete pointed metric space. For every $n\in\mathbb{N}$, let $\gamma_{n}=\sum_{k}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}\in S_{\mathcal{F}(M)}$ be such that $\sum_{k}d(x_{k}^{n},y_{k}^{n})\to 0$ and $\sup_{n}\sum_{k}\lambda_{k}^{n}<\infty$. Then $(\gamma_{n})_{n}$ has a subsequence which is an (asymptotically isometric) $\ell_{1}$-basis.
@@ 1.2. Lipschitz free spaces
Then $\mathsf{M}_{h}$ is a weak-to-weak continuous operator from $\operatorname{Lip}_{0}(K)$ into $\operatorname{Lip}_{0}(M)$, called a multiplication operator. Its preadjoint $T_{h}:\mathcal{F}(M)\to\mathcal{F}(K)$, given by $T_{h}(\delta(x))=h(x)\,\delta(f(x))$, is referred to as a weighted Lipschitz operator. Precise conditions on $h$ and $f$ ensuring boundedness of $\mathsf{M}_{h}$ and $T_{h}$ are characterized in [2].
@@ 3. $(iii)$
Let $\gamma_{1}\in W$ be such that $\|\gamma_{1}\|>\varepsilon$. Let $\mu_{1}\in\mathcal{F}(B(0,r))$ be such that $\|\gamma_{1}-\mu_{1}\|\leq\frac{\varepsilon}{4}$. Then $\|\mu\|>\frac{3\varepsilon}{4}$ and thus there exists $g_{1}\in B_{\operatorname{Lip}_{0}(M)}$ such that $\langle g_{1},\mu_{1}\rangle>\frac{3\varepsilon}{4}$. We may additionally assume that $g_{1}\mathord{\upharpoonright}_{B(0,2r)^{c}}=0$ by replacing $g_{1}$ with a function $g^{\prime}_{1}$ defined by $g^{\prime}_{1}=g_{1}$ on $B(0,r)$, $g^{\prime}_{1}=0$ on $B(0,2r)^{c}$, and extended to $M$ by McShane-Whitney theorem (see e.g. [25, Theorem 1.33]). Furthermore, $\langle g_{1},\gamma_{1}\rangle>\frac{\varepsilon}{2}$. Pick a finitely supported element $\gamma_{1}^{\prime}\in\mathcal{F}(M)$ such that $\langle g_{1},\gamma_{1}^{\prime}\rangle>\frac{\varepsilon}{2}$ and $\|\gamma_{1}-\gamma_{1}^{\prime}\|<\frac{\varepsilon\delta}{16r}$. Let $F_{1}=\operatorname{supp}(\gamma_{1}^{\prime})\cup\left\{{0}\right\}$ and define $f_{1}\colon[F_{1}]_{\delta/2}^{c}\cup F_{1}\rightarrow\mathbb{R}$ in such way that $f_{1}\mathord{\upharpoonright}_{[F_{1}]_{\delta/2}^{c}}=0$ and $f_{1}\mathord{\upharpoonright}_{F_{1}}=\frac{\delta}{4r}g_{1}\mathord{\upharpoonright}_{F_{1}}$. Since $g_{1}\mathord{\upharpoonright}_{B(0,2r)^{c}}=0$, for every $u\in F_{1}$ we have $\left|{g_{1}(u)}\right|\leq 2r$ and thus $\left|{f_{1}(u)}\right|\leq\frac{\delta}{2}$. Thus, since $\frac{\delta}{4r}\leqslant 1$, the Lipschitz constant of $f_{1}$ is at most 1. Extend $f_{1}$ to $M$ using McShane–Whitney Theorem. Then $f_{1}\in B_{\operatorname{Lip}_{0}(M)}$ and $\langle f_{1},\gamma_{1}^{\prime}\rangle=\frac{\delta}{4r}\langle g_{1},\gamma_{1}^{\prime}\rangle>\frac{\varepsilon\delta}{8r}$, and thus $\langle f_{1},\gamma_{1}\rangle>\frac{\varepsilon\delta}{16r}$.
@@ 3. Uniform regularity versus weak compactness
Note that for every $i\in\{1,\ldots,n\}\setminus I$, there exist $x_{i}\in B(u_{i},r_{i})\cap U^{c}$ and $y_{i}\in B(v_{i},r_{i})\cap U^{c}$. Thus, by [23, Lemma 1.2]:
@@
for every $k\in\mathbb{Z}$. The well-known Kalton decomposition [21, Proposition 4.3] asserts that there exist Lipschitz functions $\varphi_{n}:M\to\mathbb{R}$ with $\operatorname{supp}(\varphi_{n})\subset M_{n}$ such that the pre-adjoints $T_{n}=T_{\varphi_{n}}$ of the associated multiplication operators $\mathsf{M}_{\varphi_{n}}:\operatorname{Lip}_{0}(M_{n})\to\operatorname{Lip}_{0}(M)$ satisfy that
@@ 2. $(b)$
Notice that when $\mu$ is finite, condition (a) implies condition (b). According to the Dunford–Pettis theorem, a set $H$ is relatively weakly compact in $L^{1}(\mu)$ if and only if it is equi-integrable (see e.g. [24, Theorem 15.4]). It is worth noting that, since $L^{1}$-spaces are weakly sequentially complete (see [3, Theorem 5.2.9]), weakly precompact sets coincide with relatively weakly compact sets. In what follows, we will transfer the equi-integrability conditions through the isometry $L^{1}(\lambda_{T})\to\mathcal{F}(T)$ and describe the resulting properties in $\mathcal{F}(T)$.
@@ 5. Sequences of sums of “small molecules” contain an $\ell_{1}$-basis
The previous section highlights the significance of approximation by convex series of molecules. Informally, within a weakly compact set, the weight of molecules formed by pairs of points in close proximity should be controlled in a uniform way. In this section, we establish that this condition is, in a sense, necessary. Specifically, consider a sequence of (finite or countable) linear combinations of molecules, denoted by $\gamma_{n}=\sum_{k}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}$, satisfying $\sum_{k}d(x_{k}^{n},y_{k}^{n})\to 0$ as $n\to\infty$ and $\sup_{n}\sum_{k}|\lambda_{k}^{n}|<\infty$. We will demonstrate that this sequence contains an $\ell_{1}$-basis, thereby proving that such a sequence cannot be contained in a weakly precompact set.
We shall start by introducing a definition of $\delta$-cyclical monotonicity between subsets of $\widetilde{M}:=\left\{{(x,y):x,y\in M,x\neq y}\right\}$. For a subset of $\widetilde{M}$, the concept of cyclical monotonicity can be found e.g. in [11]: the set $A\subset\widetilde{M}$ is cyclically monotonic if, for every finite sequence of pairs $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\in A$, we have
$\sum_{i=1}^{n}d(x_{i},y_{i+1})\geqslant\sum_{i=1}^{n}d(x_{i},y_{i}),$
where $y_{n+1}=y_{1}$. This concept was generalized in [18] by introducing a parameter $\delta\in(0,1]$: a subset $A\subset\widetilde{M}$ is $\delta$-cyclically monotonic, if for any finite sequence of pairs $(x_{1},y_{1})$, $\ldots$ , $(x_{n},y_{n})\in A$, we have
$\sum_{i=1}^{n}\min\left\{{d(x_{i},y_{i+1})-\delta d(x_{i},y_{i}),d(y_{i},y_{i+1})}\right\}\geq 0,$
where $y_{n+1}:=y_{1}$. Note that $A$ is cyclically monotonic if and only if it is 1-cyclically monotonic. The motivation for both [11] and [18] to consider these definitions was to find a lower estimate for the norm of the sum of molecules. In particular, by [18, Proposition 2.2] (or by [11, Theorem 2.4] for the case $\delta=1$), a set $A\subset\widetilde{M}$ is $\delta$-cyclically monotonic if and only if there exists $f\in S_{\operatorname{Lip}_{0}(M)}$ such that $\langle f,m_{xy}\rangle\geq\delta$ for every $(x,y)\in A$.
Definition 5.1.
Let $A,B\subset\widetilde{M}$ and let $\delta\in(0,1]$. We say that the sets $A$ and $B$ are $\delta$-cross-cyclically monotonic, if for any finite sequence of pairwise distinct pairs $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\in A\cup B$, which contains at least one pair from each of $A$ and $B$, we have
(3) $\sum_{i=1}^{n}\min\left\{{d(x_{i},y_{i+1})-\delta d(x_{i},y_{i}),d(y_{i},y_{i+1})}\right\}\geq 0,$
where $y_{n+1}:=y_{1}$. When the sets $A$ and $B$ are 1-cross-cyclically monotonic, the inequality (3) reduces to
$\sum_{i=1}^{n}\left(d(x_{i},y_{i+1})-d(x_{i},y_{i})\right)\geq 0,$
and in this case, we simply say that the sets $A$ and $B$ are cross-cyclically monotonic.
Remark 5.2.
Notice that when $A$ and $B$ are $\delta$-cross-cyclically monotonic (with $\delta>0$), then $(x,y)\in A$ implies $(y,x)\notin B$. Nevertheless this does not prevent both $(x,y)$ and $(y,x)$ from belonging to $A$. Indeed, let $M=\left\{{x,y,a,b}\right\}$ such that $d(x,y)=d(a,b)=1$ and every other non-zero distance is $2$. Then $A=\left\{{(x,y),(y,x)}\right\}$ and $B=\left\{{(a,b),(b,a)}\right\}$ are cross-cyclically monotonic, since for any sequence of pairwise distinct elements $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\in A\cup B$, $n=2,3,4$, that contains elements from both sets, we have
$\sum_{i=1}^{n}d(x_{i},y_{i+1})\geq 4\geq\sum_{i=1}^{n}d(x_{i},y_{i}),$
where $y_{i+1}=y_{1}$. Note that $A$ and $B$ are clearly not $\delta$-cyclically monotonic for any $\delta>0$. In particular, cross-cyclical monotonicity places no cyclical-monotonicity constraint on the individual components.
Remark 5.3.
We need the assumption that the pairs $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ are pairwise distinct to avoid the situation when we select the same pair twice from the same set. However, it is straightforward to check that, when the sets $A$ and $B$ are cross-cyclically monotonic, then the inequality (3) holds also for all such sequences $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\in A\cup B$, where whenever two pairs $(x_{k},y_{k})$ and $(x_{k^{\prime}},y_{k^{\prime}})$ are equal, then they belong to different sets $A$ and $B$.
We now extend [11, Theorem 2.4] and [18, Proposition 2.2] to the setting of mutually $\delta$-cyclically monotonic sets.
Proposition 5.4.
Let $\delta\in(0,1]$ and let $A_{1},A_{2}\subset\widetilde{M}$ be $\delta$-cross-cyclically monotonic. Let $\lambda^{i}\in\ell_{1}^{+}(A_{i})$, and denote $\gamma_{i}=\sum_{(x,y)\in A_{i}}\lambda^{i}_{(x,y)}m_{xy}$, for $i=1,2$. Then there exists $h\in S_{\operatorname{Lip}_{0}(M)}$ such that $\left<{h,\gamma_{i}}\right>\geqslant\delta\left\|{\gamma_{i}}\right\|$ for $i=1,2$. In particular $\left\|{\gamma_{1}+\gamma_{2}}\right\|\geqslant\delta(\left\|{\gamma_{1}}\right\|+\left\|{\gamma_{2}}\right\|)$.
[…]
@@
Let $I$ be a set and $\beta_{ij}$, $i,j\in I$, be real numbers such that $\beta_{ii}=0$ for every $i\in I$. Then the following statements are equivalent:
  • (i)
    There exist numbers $\alpha_{i}$, $i\in I$ such that
    $\alpha_{i}\leqslant\alpha_{j}+\beta_{ij}$
    for $i,j\in I$.
  • (ii)
    For every finite sequence $i_{1},\ldots,i_{m}$ in $I$ we have
    $\beta_{i_{1}i_{2}}+\beta_{i_{2}i_{3}}\ldots+\beta_{i_{m}i_{1}}\geqslant 0,$
  • (iii)
    For every injective finite sequence $i_{1},\ldots,i_{m}$ in $I$ we have
    $\beta_{i_{1}i_{2}}+\beta_{i_{2}i_{3}}\ldots+\beta_{i_{m}i_{1}}\geqslant 0,$
Proof of Proposition 5.4.
The proof follows similar ideas as [11, Theorem 2.4] and [18, Proposition 2.2]. Let $f,g\in S_{\operatorname{Lip}_{0}(M)}$ be such that $\langle f,\gamma_{1}\rangle=\left\|{\gamma_{1}}\right\|$ and $\langle g,\gamma_{2}\rangle=\left\|{\gamma_{2}}\right\|$. Write $A_{1}=\left\{{(x_{i},y_{i}):i\in I_{1}}\right\}$ and $A_{2}=\left\{{(x_{i},y_{i}):i\in I_{2}}\right\}$, where $I_{1}$ and $I_{2}$ are disjoint sets of indices. Let $I=I_{1}\cup I_{2}$. For $i,j\in I$, let
$\beta_{ij}:=\begin{cases}\min\left\{{d(x_{i},y_{j})-\delta\left(f(x_{i})-f(y_{i})\right),d(y_{i},y_{j})}\right\},&\text{ if }i\in I_{1}\\ \min\left\{{d(x_{i},y_{j})-\delta\left(g(x_{i})-g(y_{i})\right),d(y_{i},y_{j})}\right\},&\text{ if }i\in I_{2}.\end{cases}$
Clearly $\beta_{jj}=0$ for all $j\in I$. Let us consider pairwise distinct indices $i_{1},\ldots,i_{n}\in I$. Our aim is to show that
$\sum_{k=1}^{n}\beta_{i_{k}i_{k+1}}\geq 0,$
where we denote $i_{n+1}:=i_{1}$. First consider the case when $i_{1},\ldots,i_{n}\in I_{1}$. Then
$\sum_{k=1}^{n}\beta_{i_{k}i_{k+1}}$ $=\sum_{k=1}^{n}\min\left\{{d(x_{i_{k}},y_{i_{k+1}})-\delta\left(f(x_{i_{k}})-f(y_{i_{k}})\right),d(y_{i_{k}},y_{i_{k+1}})}\right\}$
$\geq\delta\sum_{k=1}^{n}\min\left\{{d(x_{i_{k}},y_{i_{k+1}})-f(x_{i_{k}})+f(y_{i_{k}}),d(y_{i_{k}},y_{i_{k+1}})}\right\}$
$\geq\delta\sum_{k=1}^{n}\left(-f(y_{i_{k+1}})+f(y_{i_{k}})\right)=0.$
The case when $i_{1},\ldots,i_{n}\in I_{2}$ is analogous. Now assume that $i_{1},\ldots,i_{n}$ contain indices from both sets $I_{1}$ and $I_{2}$. For all $i,j\in I$, we have
$\beta_{ij}\geq\min\left\{{d(x_{i},y_{j})-\delta d(x_{i},y_{i}),d(y_{i},y_{j})}\right\},$
and thus we obtain $\sum_{k=1}^{n}\beta_{i_{k}i_{k+1}}\geq 0$ from the fact that the sets $A_{1}$ and $A_{2}$ are $\delta$-cross-cyclically monotonic and from Remark 5.3. Indeed, notice that if for $i,j\in I$ with $i\neq j$ we have $(x_{i},y_{i})=(x_{j},y_{j})$ then necessarily $(x_{i},y_{i})\in A_{1}$ and $(x_{j},y_{j})\in A_{2}$ or vice versa. Next, from Lemma 5.5 we obtain real numbers $\alpha_{i}$, $i\in I$ such that
$\alpha_{i}\leq\alpha_{j}+\beta_{ij}$
for all $i,j\in I$. Define $h:M\rightarrow\mathbb{R}$ by
$h(u)=\inf_{i\in I}(\alpha_{i}+d(u,y_{i})).$
As the infimum of 1-Lipschitz functions, $h$ is 1-Lipschitz. The properties of $\alpha_{i}$, $i\in I$, are not affected by adding a constant, thus we may additionally assume they were chosen in such a way that $h(0)=0$ and thus $h\in B_{\operatorname{Lip}_{0}(M)}$. It remains to show that $\langle h,\gamma_{i}\rangle\geq\delta\left\|{\gamma_{i}}\right\|$, for $i=1,2$. Fix $i\in I_{1}$ and $\varepsilon>0$. There exists $j\in I$ such that $h(x_{i})\geq\alpha_{j}+d(x_{i},y_{j})-\varepsilon$. Furthermore, $h(y_{i})\leq\alpha_{i}$, and thus
$h(x_{i})-h(y_{i})\geq\alpha_{j}+d(x_{i},y_{j})-\varepsilon-\alpha_{i}\geq d(x_{i},y_{j})-\beta_{ij}-\varepsilon\geq\delta\left(f(x_{i})-f(y_{i})\right)-\varepsilon.$
As $\varepsilon>0$ was arbitrary, we obtain $h(x_{i})-h(y_{i})\geq\delta\left(f(x_{i})-f(y_{i})\right)$ for every $i\in I_{1}$, and thus $\langle h,\gamma_{1}\rangle\geq\delta\langle f,\gamma_{1}\rangle=\delta\left\|{\gamma_{1}}\right\|$. By the same argument, we obtain $h(x_{i})-h(y_{i})\geq\delta\left(g(x_{i})-g(y_{i})\right)$ for every $i\in I_{2}$ and thus $\langle h,\gamma_{2}\rangle\geq\delta\langle g,\gamma_{2}\rangle=\delta\left\|{\gamma_{2}}\right\|$. This concludes the proof. ∎
Lemma 5.6.
Let $M$ be a pointed metric space. Let $\mu:=\sum_{k}\lambda_{k}m_{x_{k}y_{k}}\in\mathcal{F}(M)$ where $(\lambda_{k})\in\ell_{1}$ and let $\gamma\in\mathcal{F}(M)$. Let $\delta:=\sum_{k}d(x_{k},y_{k})$ and
@@
By an obvious approximation argument, we may assume that $\gamma$ and $\mu$ are finitely supported, with $\mu$ written as $\mu:=\sum_{k=1}^{n}\lambda_{k}m_{x_{k}y_{k}}$. We may additionally assume $\lambda_{k}>0$ by interchanging $x_{k}$ and $y_{k}$ whenever necessary. Let $A=\operatorname{supp}(\gamma)\cup\left\{{0}\right\}$ and $B=\left\{{x_{1},y_{1},\ldots,x_{n},y_{n}}\right\}$. We may further assume that $\gamma$ is written as a convex sum of molecules from points in $A$, say $m_{x_{n+1}y_{n+1}},\ldots,m_{x_{N}y_{N}}$, where $x_{n+1},y_{n+1},\ldots,x_{N},y_{N}\in A$. Finally, we assume without loss of generality that $M=A\cup B$. Note that $A\cap B=\emptyset$ as $d(A,B)=C>\delta>0$.
@@
Claim: The sets $\left\{{(x_{1},y_{1}),\ldots,(x_{n},y_{n})}\right\}$ and $\left\{{(x_{n+1},y_{n+1}),\ldots,(x_{N},y_{N})}\right\}$ are cross-cyclically monotonic in the metric $b$.
Let us assume that the claim is proved. By Proposition 5.4 we get
$\left\|{\mu+\gamma}\right\|_{\mathcal{F}(M,b)}=\left\|{\mu}\right\|_{\mathcal{F}(M,b)}+\left\|{\gamma}\right\|_{\mathcal{F}(M,b)}.$
Furthermore, by definition of $b$, we have $\left\|{\mu}\right\|_{\mathcal{F}(M,b)}=\left\|{\mu}\right\|_{\mathcal{F}(M,d)}$ and $\left\|{\gamma}\right\|_{\mathcal{F}(M,b)}=\left\|{\gamma}\right\|_{\mathcal{F}(M,d)}$, and thus
$\|\mu\|_{\mathcal{F}(M,d)}+\|\gamma\|_{\mathcal{F}(M,d)}\leq\left\|{\mu+\gamma}\right\|_{\mathcal{F}(M,b)}\leq\left({1+\frac{\delta}{C}}\right)\left\|{\mu+\gamma}\right\|_{\mathcal{F}(M,d)}.$
To conclude the proof, it only remains to prove the claim. Note that, since $\gamma$ is a convex sum of the molecules $m_{x_{n+1}y_{n+1}},\dots,m_{x_{N}y_{N}}$, by [11, Theorem 2.4] the sequence $(x_{k},y_{k})_{k=n+1}^{N}$ is cyclically monotonic with respect to $d$, hence also with respect to $b$. Now let us consider $i_{1},\ldots,i_{m}\in\{1,\ldots,N\}$, with $m\geqslant 2$ and indices from both sets $\{1,\ldots,n\}$ and $\{n+1,\ldots,N\}$. Note that we may assume that the indices $i_{1},\ldots,i_{m}$ are pairwise distinct. Let $j_{1}<\ldots<j_{l}$ be such that the indices $i_{j_{k}}$, $1\leqslant k\leqslant l$, are those which belong to the set $\left\{{n+1,\ldots,N}\right\}$, and let us write $I=\{j_{1},\ldots,j_{l}\}$. To simplify the notation in the next estimates, we let $j_{l+1}:=j_{1}$, $i_{m+1}:=i_{1}$. Using the cyclic monotonicity of $(x_{k},y_{k})_{k=n+1}^{N}$ and the definition of $\delta$, we first obtain:
@@
Let $M$ be a pointed metric space. Let $\mu:=\sum_{k}\lambda_{k}m_{x_{k}y_{k}}\in\mathcal{F}(M)$ where $(\lambda_{k})\in\ell_{1}$ and let $\gamma\in\mathcal{F}(M)$. Let $r>0$ be such that for every $k$ there exists $u\in\operatorname{supp}(\gamma)\cup\left\{{0}\right\}$ such that
@@
Using the two previous lemmas, we obtain the following generalization of [20, Theorem 2.6], which covers the case where each $\gamma_{n}$ is a molecule.
@@
Let $(\gamma_{n})$ be a sequence in $S_{\mathcal{F}({M})}$ such that each $\gamma_{n}$ is expressed as follows:
$\gamma_{n}=\sum_{k}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}.$
If $\lim_{n\to\infty}\sum_{k}d(x_{k}^{n},y_{k}^{n})=0,$ and $\sup_{n}\sum_{k}\left|{\lambda_{k}^{n}}\right|<\infty$, then $\lim_{n}\left\|{\gamma+\gamma_{n}}\right\|=\left\|{\gamma}\right\|+1$ for all $\gamma\in\mathcal{F}({M})$.